Question Guestion 10 points One promanent physaan darns the 80% of those with kung cancer are chain smokers # this assertion is correct probability that of 20 of such patents recenly admited to a hospital 13 are chain smokers O a 0 00000017 O b 013 O c 0175 O d 0055
Answer #1

Solution:

The correct option is D

0.055

Mean μ = n-p = (20) . = 16. (20) . ( :) (-(:))부-a2. variance σ2-n . p. (1-p) 53.2 Standard deviation 4))-4V6 1 -p 一 1.7888543819998 P(X 1 3) 0.0545498504865251.